OpenStudy (idealist10):

How to integrate (4-4x)/(1-2x) dx?

OpenStudy (idealist10):

\[\frac{ 4-4x }{ 1-2x }dx\]

myininaya (myininaya):

\[\frac{4-4x}{1-2x}=\frac{2(1-2x)+2}{1-2x}=2+\frac{2}{1-2x}\]

myininaya (myininaya):

do you know how to integrate each of those terms?

OpenStudy (idealist10):

Yes, but how did you get \[2+\frac{ 2 }{ 1-2x }\]

myininaya (myininaya):

my middle step there

myininaya (myininaya):

4-4x=2(1-2x)+2

myininaya (myininaya):

then you just separate the fraction

OpenStudy (idealist10):

Oh! Now I got it! How come I didn't thought of that before? :) Thank you!

myininaya (myininaya):

you can also just use straight up long division

OpenStudy (idealist10):

You're very helpful!

myininaya (myininaya):

lol thanks